X Company makes two products, A and B, and uses an activity-based costing overhead allocation system, with three cost pools and three cost drivers. Budgeted costs and driver information for 2017 were as follows:

Cost Pool 1 Cost Pool 2 Cost Pool 3
Budgeted costs $113,000     $58,000     $26,000    
Cost Drivers
  Product A 78,000     62,000     73,000    
  Product B 67,000     54,000     55,000    


What was the allocation to Product B in 2017?
Answer #1

Answer:

Budgeted Activity

Activity Cost Pool

Product A

Product B

Total
Activity

A

B

C=A+B

Cost pool 1

78,000

67,000

145,000

Cost pool 2

62,000

54,000

116,000

Cost pool 3

73,000

55,000

128,000

Total Cost

Total
Activity

Rate

A

B

C=A/B

Cost pool 1

113000

145,000

0.78

Cost pool 2

58000

116,000

0.50

Cost pool 3

26000

128,000

0.20

Expected

Total

Rate

Uses of cost
driver

Activity

B

C=A*B

Cost pool 1

0.78

67000

52213.8

Cost pool 2

0.50

54000

27000.0

Cost pool 3

0.20

55000

11171.9

Allocation to Product B

90385.7

Allocation to Product B

90385.7